# Natural polynomials for Kerr quasi-normal modes

### v2

@@ VI Tridiagonalization
In Eq. (VI), as in Eq. (11), $\mathcal{L}_{\xi}$ has been written in terms of the physical constants $\text{C}_{0}$ and $\text{C}_{1}$ (Eq. 62a), as well as the confluent Heun operator $\mathcal{D}_{\xi}$ (Eq. 62b). Plainly, Eq. (62a) only contains constant rescalings of $|{u_{p}}\rangle$ and terms dependent on $\xi|{u_{q}}\rangle$, and Eq. (62b) is simply $\mathcal{D}_{\xi}\,|{u_{q}}\rangle$. Since ${\langle{u_{p}}\,|\,{u_{q}}\rangle}=\delta_{pq}$, constant rescalings of $|{u_{p}}\rangle$ will only contribute to the diagonal of $\hat{L}$. Therefore, to understand the structure of $\hat{L}$, we must first refine our understanding of $\xi|{u_{q}}\rangle$ and $\mathcal{D}_{\xi}\,|{u_{q}}\rangle$.
@@
In the context of |$\mathcal{L}{{}_{\xi}}\,|{u_{q}}\rangle$ (i.e. Eq. VI), the existence of a three-term recursion relationship between polynomials means that terms within Eq. (62a) only contribute to the central, lower and upper diagonal of $\hat{L}$. Therefore, Eq. (62a) will generally contribute to the tridiagonal part of $\hat{L}$.
@@
Further, since $\mathcal{D}_{\xi}$ contains derivatives, it is here that the representation of each $|{u_{q}}\rangle$ has meaningful effect: while the previously discussed single- and multi-order representations are equivalent (Eqs. V and V), their behaviors under differentiation are not.
@@
It follows (from Eqs. VI and 66c) that $\mathcal{L}_{\xi}\,|{u_{q}}\rangle$ may be written as
@@
In Eq. (68c), recall that $\alpha_{p^{\prime}}$ is defined in Eq. (36). The structure of Eq. (VI) enables a few useful observations. Equations (68a-68c) imply that $\hat{L}$ is the sum of three matrices, $\hat{L}^{(\text{S})}_{0}$, $\hat{L}^{(\text{S})}_{1}$, and $\hat{L}^{(\text{S})}_{2}$,
@@
Note that ${\langle{u_{q}}\,|\,{v_{p^{\prime}p}}\rangle}={\langle{v_{p^{\prime}p}}\,|\,{u_{q}}\rangle}$, not due to Eq. (71), but rather to the multiplicative associative property applied to the integrand of the scalar product (Eq. 14). Also note that the conditional in Eq. (73b) is equivalent to $p<q-1$.
@@
In Eq. (VI), we recall from Eq. (36) that $\alpha_{p^{\prime}}=\mu_{p^{\prime}}-\text{C}_{1}$, and that $\sigma_{q0}$, $\sigma_{q1}$ and $\sigma_{q2}$ are defined by Eq. (VI).
@@
Plainly stated, since $\hat{L}$ is a symmetric matrix its matrix elements, $L_{pq}$, are such that $L_{pq}=L_{qp}$. Equation (VI) means that $L_{pq}$ will be within the set $\{\sigma_{q5},\sigma_{q6},\sigma_{q7}\}$ for $p\in\{q-1,q,q+1\}$, respectively. Since $\hat{L}$ is tridiagonal, the only nontrivial symmetry statement is $L_{q-1,q}=L_{q,q-1}$, which is equivalent to $L_{q,q+1}=L_{q+1,q}$ under $q\rightarrow q+1$. Using equation Eqs. (84) and (VI), $L_{q-1,q}=L_{q,q-1}$ becomes,
@@
In Eq. (86a), $L_{q-1,q}=L_{q,q-1}$ has been rewritten using Eq. (84). In going from Eq. (86a) to Eq. (86b), Eq. (VI) has been used to expand both sides. Equation (86b) may be further reduced by noting that symmetry of the canonical polynomials’ three-term recursion (See Eq. 63) means that $\sigma_{q-1,2}=\sigma_{q0}$. Therefore, Eq. (86b) allows $\sigma_{q3}$ (i.e. ${\langle{u_{q-1}}\,|\,{v_{q}}\rangle}$) to be written as,
@@ VII Select numerical results
Before being put into practice, the preceding theoretical development benefits from a few points of guidance London (2023a). For all computations, physical parameters, namely $\{a/M,s,\tilde{\omega},m\}$ must be chosen such that the derived parameters, ${\text{C}_{1}}$ through ${\text{C}_{4}}$, may then be computed according to Eq. (III). For nearly all subsequent manipulations involving polynomials or series expansions, the fundamental operation is the scalar product (See Eq. 14).
@@
As seen in the top panel of Fig. 2, the polynomials may be formatted such that their real parts have zero roots along the real line, meaning that the net complex valued polynomials will have no roots at real values of the fractional radial coordinate, $\xi$. This is simply a result of the fact that the canonical confluent Heun polynomials have complex valued coefficients (Eq. V), and so will generally have roots situated away from the real line.
@@
–  The canonical confluent Heun polynomials may be used to tridiagonalize confluent Heun equations such as the QNM’s radial equation (Eq. III), or the confluent Heun polynomial equation (Eq. III).
@@
Conceptually, canonical confluent Heun polynomial orthogonality (Eq. 41) is demonstrated in panel $(a)$, upper-tridiagonality of auxiliary polynomials (Eq. 82) is demonstrated in panel $(c)$, and three-term recursion (Eq. VI) is dempstrated in panel $(e)$. Then end result, panel $(g)$, may be considered to represent the three-term recursion of orthogonal polynomials which approximate solutions to the QNM radial problem Teschl (2000); Kuijlaars et al. (2003).
@@
Since $k$ and $\alpha$ are both integers between 00 and $p$, the sequence of all eigenvectors comprise the rows of a matrix whose elements are ${y}_{p\alpha,\beta}$. Similarly, the eigenvalues may be labeled as $\lambda_{p\alpha}$ rather than $\lambda_{pk}$.
@@
Figure 5: Radial eigenvalue distributions for select quasinormal mode cases, all with BH dimensionless spin $a/M=0.7$, and angular indices $(\ell,m)=(2,2)$. In all panels, radial eigenvalues, $A_{n^{\prime}}$, are noted with open circles, and the known QNM eigenvalue, $A_{\ell mn}$, is shown as a red filled circle (c.f. Fig. 1). Left to right: Radial eigenvalues for the respective $n\in\{0,3,6,12\}$ QNM overtones. Respective absolute differences between known QNM eigenvalues (Refs. London et al. (2020); Stein (2019)) and those computed here (i.e. $|A_{\ell mn}-A_{n}|$) are $\{2.72\,\times 10^{-9},4.72\,\times 10^{-9},1.24\,\times 10^{-9},6.28\,\times 10^{-9}\}$. Center right ($n=6$): Eigenvalues for the closest confluent Heun polynomial problem are shown as black triangles. See panel $(c)$ of Fig. 4 for comparison.
[…]
This is consistent with the so-called “$\Delta_{p+1}$ condition” which is used to define the confluent Heun polynomials via a finite dimensional matrix eigenvalue problem (see Eq. 79 of Paper I as well as Refs. DLMF ; Cook and Zalutskiy (2014)). There the basic concept is that a power series solution to Eq. (89b) will terminate after $p+1$ terms if the $\Delta_{p+1}$ condition on a related three-term recursion is satisfied.
The $\Delta_{p+1}$ condition is unable to provide insight on the non-polynomial solutions, simply due to fact that these solutions have series expansions of infinite length. In other words, the non-polynomial solutions satisfy an infinite dimensional matrix eigenvalue problem, rather than a finite dimensional one555 In such cases, standard routes to finding solutions include the use of continued fractions (see e.g. Ref. Leaver (1985)) or the representation of the problem with an orthonormal basis, such as the canonical confluent Heun polynomial basis introduced here. . Unlike the power-series perspective of the $\Delta_{p+1}$ condition, the canonical confluent Heun polynomial representation of the problem naturally applies to finite and infinite dimensional problems: i.e., the monomials are the basis for power series expansions, but unlike the canonical confluent Heun polynomial, monomials are not an orthonormal basis, meaning that standard matrix interpretations do not apply for non-terminating series. Since the canonical confluent Heun polynomial are orthonormal, the polynomial and non-polynomial sectors are clearly visible in Fig. 4a.
@@
Since the imaginary part of ${p_{\star}^{+}}$ is small, and the real part is close to $6$ (c.f. Eq. IV), it is reasonable to expect rough agreement between that QNM’s radial eigenvalues, and the order-$6$ confluent Heun polynomial’s polynomial and non-polynomial eigenvalues (i.e. ${p_{\star}^{+}}\approx 6$).
In Fig. 5, the radial eigenvalue distributions for four select QNMs are shown. There, as $n$ increases, so does the approximately polynomial sector of the eigenvalue distribution. For the $n=6$ overtone, the order-$6$ confluent Heun polynomial eigenvalues have been added for comparison. In that case, it is clear that the qualitative features of the QNM eigenvalue distribution are precisely captured by that of the confluent Heun polynomial eigenvalues. Further, it is clear that, for this case, the QNM separation constant, $A_{226}$, is indeed well approximated by the dominant order-$6$ confluent Heun polynomial eigenvalue,
@@
In Eq. (94), Eq. (31) has been restated in the current context, and the parameters, ${\text{C}_{0}}$ and ${\text{C}_{1}}$, have been computed with the relevant QNM frequency $\tilde{\omega}_{226}$. The fractional error between the left and right hand sides of Eq. (94) is $4.6\%$.
@@
In Eq. (98), Eq. (95) has been rearranged such that all quantities appear on the left hand side. The absolute value has then been taken to yield a single, real-valued error measure.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Radial functions, $f_{n}(\xi)$, and related floating point error estimates, $\Delta_{n}$, for the same cases shown in Fig. 5. Top row: The real part of each radial function, $\text{Re}\,f_{n}(\xi)$, is denoted with a solid orange line. The imaginary part, $\text{Im}\,f_{n}(\xi)$, is denoted with a dashed purple line, and the function’s positive and negative magnitude, ±$\pm|f_{n}(\xi)|$ are denoted with dotted black lines. Bottom row: errors for each radial function, $\Delta_{n}$ (Eq. 98). See Sec. VII for related discussion.
If $f_{n}(\xi)$ are eigenfunctions, and the derivatives within $\mathcal{L}_{\xi}$ can be evaluated with infinite precision, then $\Delta_{n}(\xi)$ would equal zero for all values of $\xi$. In practice, $\Delta_{n}(\xi)$ will be very small relative to one, and dominated by numerical noise.
@@
Further, due to the self-adjoint nature of the radial operator with respect to the scalar product (Eq. III), the basis of radial functions is orthogonal London (2023a). Since the canonical confluent Heun polynomial are an orthonormal basis, scalar products between different radial functions, $|{f_{p}}\rangle$ and $|{f_{q}}\rangle$, are equivalent to the simple vector dot product between their vector representations,
@@
Consequently, the radial functions may be normalized in the standard way, $|{f_{p}}\rangle\leftarrow|{f_{p}}\rangle/\sqrt{{\langle{f_{p}}\,|\,{f_{p}}\rangle}}$. In other words, when computed for a single frequency parameter, the radial functions define an orthonormal basis.
A central question in single BH perturbation theory is whether QNMs for one physical case may be used to efficiently represent QNMs in another. For example, it is extremely common to represent the Kerr QNMs’ angular functions in the basis of those for Schwarzschild London (2023a, b); Berti and Klein (2014); O’Sullivan and Hughes (2014); Cook and Zalutskiy (2014); Press and Teukolsky (1973). Here, we are concerned with whether the same is true for the QNMs’ radial functions. In effect, we ask whether there exists a mixing matrix between sets of radial functions derived from different BH spins.
To answer this question, it cannot be avoided that, even for a single QNM, different values of BH spin correspond to different QNM frequencies London and Fauchon-Jones (2019); Leaver (1985); Berti and Klein (2014); Berti et al. (2016). When combined with the fact that the scalar product’s weight function (Eq. 16) is frequency dependent, this means that the QNM radial functions for different BH spins are members of different scalar product spaces. Therefore, a complete and selfconsistent treatment of the question requires the use of a frequency-independent weight function. While such a weight function was defined in Sec. IV-B of Paper I, we consider a complete treatment of its use to be outside of the current scope.
In lieu of a more rigorous treatment, Eq. (VII) may be used to construct a simplified one. Let the $p^{\text{th}}$ radial function for a given BH spin, $a$, be denoted $\vec{f}_{p}^{\,(a)}$. As mentioned previously, if $\vec{f}_{p}^{\,(a)}$ are defined with QNM frequency $\tilde{\omega}_{{\ell mn}}$, then $\vec{f}_{n}^{\,(a)}$ is the QNM’s radial function. Since each $a$-dependent canonical confluent Heun polynomial basis is orthonormal, Eq. (VII) may be interpreted to mean that the scalar product for all $\vec{f}_{p}^{\,(a)}$ is given by the right hand side of Eq. (99c). Since the right hand side of Eq. (99c) has no weight function, we may superficially conclude that the standard vector dot product is a frequency-independent scalar product of potential use. This conclusion is only superficial because it ignores the fact that a frequency dependent weight is still required to arrive at Eq. (99c). Nevertheless, use of Eq. (99c) for radial functions of different BH spins is sufficient for the present example.
@@
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 7: Mixing matrices between Schwarzschild and Kerr radial functions for four values of BH spin: $a/M\in\{0,0.01,0.70,0.94\}$. For each case, the first $24$ radial functions have been computed using canonical confluent Heun polynomials derived from the $s=-2$, $(\ell,m,n)=(2,2,0)$ QNM frequencies. Mixing matrices have been computed according to Eq. (100).

### v3

@@ VI Tridiagonalization
In Eq. (62), as in Eq. (11), $\mathcal{L}_{\xi}$ has been written in terms of the physical constants $\text{C}_{0}$ and $\text{C}_{1}$ (Eq. 62a), as well as the confluent Heun operator $\mathcal{D}_{\xi}$ (Eq. 62b). Plainly, Eq. (62a) only contains constant rescalings of $|{u_{p}}\rangle$ and terms dependent on $\xi|{u_{q}}\rangle$, and Eq. (62b) is simply $\mathcal{D}_{\xi}\,|{u_{q}}\rangle$. Since ${\langle{u_{p}}\,|\,{u_{q}}\rangle}=\delta_{pq}$, constant rescalings of $|{u_{p}}\rangle$ will only contribute to the diagonal of $\hat{L}$. Therefore, to understand the structure of $\hat{L}$, we must first refine our understanding of $\xi|{u_{q}}\rangle$ and $\mathcal{D}_{\xi}\,|{u_{q}}\rangle$.
@@
In the context of $\mathcal{L}{{}_{\xi}}\,|{u_{q}}\rangle$ (i.e. Eq. 62), the existence of a three-term recursion relationship between polynomials means that terms within Eq. (62a) only contribute to the central, lower and upper diagonal of $\hat{L}$. Therefore, Eq. (62a) will generally contribute to the tridiagonal part of $\hat{L}$.
@@
Further, since $\mathcal{D}_{\xi}$ contains derivatives, it is here that the representation of each $|{u_{q}}\rangle$ has meaningful effect: while the previously discussed single- and multi-order representations are equivalent (Eqs. 53 and 54), their behaviors under differentiation are not.
@@
It follows (from Eqs. 62 and 66c) that $\mathcal{L}_{\xi}\,|{u_{q}}\rangle$ may be written as
@@
In Eq. (68c), recall that $\alpha_{p^{\prime}}$ is defined in Eq. (36). The structure of Eq. (68) enables a few useful observations. Equations (68a-68c) imply that $\hat{L}$ is the sum of three matrices, $\hat{L}^{(\text{S})}_{0}$, $\hat{L}^{(\text{S})}_{1}$, and $\hat{L}^{(\text{S})}_{2}$,
@@
Note that ${\langle{u_{q}}\,|\,{v_{p^{\prime}p}}\rangle}={\langle{v_{p^{\prime}p}}\,|\,{u_{q}}\rangle}$, not due to Eq. (71), but rather to the multiplicative associative property applied to the integrand of the scalar product (Eq. 14). Also note that the conditional in Eq. (73b) is equivalent to p<q1p<q-1.
@@
In Eq. (77), we recall from Eq. (36) that $\alpha_{p^{\prime}}=\mu_{p^{\prime}}-\text{C}_{1}$, and that $\sigma_{q0}$, $\sigma_{q1}$ and $\sigma_{q2}$ are defined by Eq. (64).
@@
Plainly stated, since $\hat{L}$ is a symmetric matrix its matrix elements, $L_{pq}$, are such that $L_{pq}=L_{qp}$. Equation (85) means that $L_{pq}$ will be within the set $\{\sigma_{q5},\sigma_{q6},\sigma_{q7}\}$ for $p\in\{q-1,q,q+1\}$, respectively. Since $\hat{L}$ is tridiagonal, the only nontrivial symmetry statement is $L_{q-1,q}=L_{q,q-1}$, which is equivalent to $L_{q,q+1}=L_{q+1,q}$ under $q\rightarrow q+1$. Using equation Eqs. (84) and (85), $L_{q-1,q}=L_{q,q-1}$ becomes,
@@
In going from Eq. (86a) to Eq. (86b), Eq. (85) has been used to expand both sides. Equation (86b) may be further reduced by noting that symmetry of the canonical polynomials’ three-term recursion (See Eq. 63) means that $\sigma_{q-1,2}=\sigma_{q0}$. Therefore, Eq. (86b) allows $\sigma_{q3}$ (i.e. $\langle{u_{q-1}}\,|\,{v_{q}}\rangle$}) to be written as,
@@ VII Select numerical results
Before being put into practice, the preceding theoretical development benefits from a few points of guidance London (2023a). For all computations, physical parameters, namely $\{a/M,s,\tilde{\omega},m\}$ must be chosen such that the derived parameters, $\text{C}_{1}$} through $\text{C}_{4}$}, may then be computed according to Eq. (12). For nearly all subsequent manipulations involving polynomials or series expansions, the fundamental operation is the scalar product (See Eq. 14).
@@
As seen in the top panel of Fig. 2, the polynomials may be formatted such that their real parts have zero roots along the real line, meaning that the net complex valued polynomials will have no roots at real values of the fractional radial coordinate, $\xi$. This is simply a result of the fact that the canonical confluent Heun polynomials have complex valued coefficients (Eq. 51), and so will generally have roots situated away from the real line.
@@
–  The canonical confluent Heun polynomials may be used to tridiagonalize confluent Heun equations such as the QNM’s radial equation (Eq. 13), or the confluent Heun polynomial equation (Eq. 22).
@@
Conceptually, canonical confluent Heun polynomial orthogonality (Eq. 41) is demonstrated in panel $(a)$, upper-tridiagonality of auxiliary polynomials (Eq. 82) is demonstrated in panel $(c)$, and three-term recursion (Eq. 64) is dempstrated in panel $(e)$. Then end result, panel $(g)$, may be considered to represent the three-term recursion of orthogonal polynomials which approximate solutions to the QNM radial problem Teschl (2000); Kuijlaars et al. (2003).
Matrix asymptotics & truncation
–  Even if the operator $\mathcal{L}_{\xi}$ is self-adjoint with respect to a given scalar product, this does not guarantee that its matrix representation will have eigenvalues that accurately approximate those of $\mathcal{L}_{\xi}$. The extent of agreement between the two sets of eigenvalues depends on the asymptotic behavior of the matrix’s off-diagonal elements Petropoulou and Velázquez (2014). By asymptotic, we refer to the behavior of matrix elements, $L_{pq}$ with $q\in\{p-1,p,p+1\}$, as $p$ tends towards infinity.
For example, it is well known that the spherical harmonics are an excellent basis in which to represent the QNMs’ spheroidal harmonics O’Sullivan and Hughes (2014); Teukolsky (1973); Cook and Zalutskiy (2014); Fackerell and Crossman (2008); London (2023b). In part, this is because, when represented in a spherical harmonic basis, the spheroidal harmonic differential operator is asymptotically diagonal [^8] This can be seen computing the spherical-spheroidal mixing coefficients, and then considering the limit as $\ell\rightarrow\infty$ London (2023b); Berti and Klein (2014)., meaning that large-$\ell$ spheroidal harmonics are asymptotically equivalent spherical harmonics of the same $\ell$ London (2023b); Fackerell and Crossman (2008); Cook and Zalutskiy (2014). As a result, one can typically estimate spheroidal harmonic eigenvalues with little concern for truncation error by first casting the problem into a spherical harmonic basis, and then solving the corresponding matrix eigenvalue problem Cook and Zalutskiy (2014); Fackerell and Crossman (2008); London (2023b); O’Sullivan and Hughes (2014).
We now discuss why this is typically not the case for the QNMs’ radial problem. In short, off-diagonal matrix elements for the radial problem typically grow at least as fast as the diagonal ones, meaning that the matrix does not become asymptotically diagonal. This behavior is explicit in the radial functions’ Frobenius-type series solution as studied by Refs. Leaver (1985); Nollert (1999); Leaver (1986); Cook and Zalutskiy (2014), and is therefore necessarily present in any orthonormal polynomial representation of the same problem [^9] This is due to the fact that the 3-term recursions for the series solution must be asymptotically equivalent to the effective tridiagonalization of any polynomial representation. While many polynomial representations will not result in tridiagonal operators, all such operators can be numerically tridiagonalized using standard techniques Virtanen et al. (2020). Chung et al. (2023b); Jaramillo et al. (2021); Dias and Reall (2013). In effect, the slow convergence of the radial functions’ series solutions, along with truncation thereof, results in what is known as spectral pollution Davies and Plum (2003); Lewin and Séré (2008); Boulton (2016); Martínez Adame (2007).
For the canonical confluent Heun polynomials, panel $(e)$ of Fig. 3 illustrates a general observation, namely that the large-$p$ structure of 3-term recursion is that of a tridiagonal Toeplitz matrix (i.e. a symmetric matrix with constant elements along its diagonals). It follows from the role of $\alpha_{p}\;=\;(p-{p_{\star}^{+}})(p-{p_{\star}^{-}})$ in Eq. (85) that all $L_{pq}$ (with $q\in\{p-1,p,p+1\}$) are asymptotically equivalent to $p^{2}$. Therefore, the matrix representation of the radial operator is not asymptotically diagonal, and not self-adjoint in the sense defined in Ref. Petropoulou and Velázquez (2014). The related qualitative behavior is seen in panel $(g)$ of Fig. 3, which shows matrix elements having a local minimum near $p=q=6$, before becoming increasingly comparable as $p$ and $q$ increase.
In forthcoming numerical results, we show both the eigenvalues that result from spectral pollution, as well as those consistent with the non-truncated (infinite dimensional) eigenvalue problem. In the latter case, eigenvalues were determined by using an $N=60$ dimensional canonical confluent Heun polynomial basis, and then truncation-induced eigenvectors were removed. Consistency with the non-truncated problem was checked by comparing against the continued-fraction result [^10] The continued-fraction result is also prone to spectral pollution. However, it can typically be avoided in that context, if $\mathcal{O}(10^{4})$ terms are considered in the series solution Cook and Zalutskiy (2014); Nollert (1999). Leaver (1985); Nollert (1999).
@@
Since $k$ and $\alpha$ are both integers between 0 and $p$, the sequence of all eigenvectors comprise the rows of a matrix whose elements are ${y}_{p\alpha,\beta}$. Similarly, the eigenvalues may be labeled as $\lambda_{p\alpha}$ rather than $\lambda_{pk}$.
@@
${L}^{(p)}_{6,7}\;=\;{L}^{(p)}_{7,6}\;=\;0\;.$ (92)
This is consistent with the so-called $\Delta_{p+1}$ condition” which is used to define the confluent Heun polynomials via a finite dimensional matrix eigenvalue problem (see Eq. 79 of Paper I as well as Refs. DLMF ; Cook and Zalutskiy (2014)). There the basic concept is that a power series solution to Eq. (89b) will terminate after $p+1$ terms if the $\Delta_{p+1}$ condition on a related three-term recursion is satisfied.
The $\Delta_{p+1}$ condition is unable to provide insight on the non-polynomial solutions, simply due to fact that these solutions have series expansions of infinite length. In other words, the non-polynomial solutions satisfy an infinite dimensional matrix eigenvalue problem, rather than a finite dimensional one [^11] In such cases, standard routes to finding solutions include the use of continued fractions (see e.g. Ref. Leaver (1985)) or the representation of the problem with an orthonormal basis, such as the canonical confluent Heun polynomial basis introduced here. .
@@
Figure 5: Radial eigenvalue distributions for select quasinormal mode cases, all with BH dimensionless spin $a/M=0.7$, and angular indices $(\ell,m)=(2,2)$. In all panels, radial eigenvalues, $A_{n^{\prime}}$, are noted with open circles, and the known QNM eigenvalue, $A_{\ell mn}$, is shown as a red filled circle (c.f. Fig. 1). Left to right: Radial eigenvalues for the respective $n\in\{0,3,6,12\}$ QNM overtones. Respective absolute differences between known QNM eigenvalues (Refs. London et al. (2020); Stein (2019)) and those in open circles (i.e. $|A_{\ell mn}-A_{n}|$) are $\{2.72\,\times 10^{-9},4.72\,\times 10^{-9},1.24\,\times 10^{-9},6.28\,\times 10^{-9}\}$. Eigenvalues for the confluent Heun polynomials with order $\text{Re}\;{p_{\star}^{+}}]$ (Eq. 37) are shown by black triangles. Center right ($n=6$): See panel $(c)$ of Fig. 4 for comparison.
Unlike the power-series perspective of the $\Delta_{p+1}$ condition, the canonical confluent Heun polynomial representation of the problem naturally applies to finite and infinite dimensional problems: i.e., the monomials are the basis for power series expansions, but unlike the canonical confluent Heun polynomial, monomials are not an orthonormal basis, meaning that standard matrix interpretations do not apply for non-terminating series. Since the canonical confluent Heun polynomial are orthonormal, the polynomial and non-polynomial sectors are clearly visible in Fig. 4a.
@@
Since the imaginary part of $p_{\star}^{+}$} is small, and the real part is close to $6$ (c.f. Eq. 37), it is reasonable to expect rough agreement between that QNM’s radial eigenvalues, and the order-$6$ confluent Heun polynomial’s polynomial and non-polynomial eigenvalues (i.e. ${p_{\star}^{+}}\approx 6$).
In Fig. 5, the radial eigenvalue distributions for four select QNMs are shown. There, as $n$ increases, so does the approximately polynomial sector of the eigenvalue distribution. The nearest confluent Heun polynomial eigenvalues have been added for comparison (See Eq. 37). For all cases, the qualitative features of the QNM eigenvalue distribution are captured by that of the confluent Heun polynomial eigenvalues. For example, the $n=6$ QNM separation constant, $A_{226}$, is indeed well approximated by the dominant order-$6$ confluent Heun polynomial eigenvalue,
@@
In Eq. (94), Eq. (31) has been restated in the current context, and the parameters, $\text{C}_{0}$} and $\text{C}_{1}$}, have been computed with the relevant QNM frequency $\tilde{\omega}_{226}$. The fractional error between the left and right hand sides of Eq. (94) is $4.6\%$. A more thorough explorarion of the QNM solution space is needed to determine whether the qualitative agreement in Fig. 5 is robust.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Radial functions, $f_{n}(\xi)$, and related floating point error estimates, $\Delta_{n}$, for the same cases shown in Fig. 5. Top row: The real part of each radial function, $\text{Re}\,f_{n}(\xi)$, is denoted with a solid orange line. The imaginary part, $\text{Im}\,f_{n}(\xi)$, is denoted with a dashed purple line, and the function’s positive and negative magnitude, $\pm|f_{n}(\xi)|$ are denoted with dotted black lines. Bottom row: errors for each radial function, $\Delta_{n}$ (Eq. 98). In black are errors for $N=24$, where all eigenvectors are consistent with the infinite dimensional problem. In magenta are errors for $N=24$, where eigenvectors due to truncation are included. See Sec. VII for related discussion.
@@
In Eq. (98), Eq. (95) has been rearranged such that all quantities appear on the left hand side. The absolute value has then been taken to yield a single, real-valued error measure. If $f_{n}(\xi)$ are eigenfunctions, and the derivatives within $\mathcal{L}_{\xi}$ can be evaluated with infinite precision, then $\Delta_{n}(\xi)$ would equal zero for all values of $\xi$. In practice, $\Delta_{n}(\xi)$ will be very small relative to one, and dominated by numerical noise.
@@
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 7: Mixing matrices between Schwarzschild and Kerr radial functions for four values of BH spin: $a/M\in\{0,0.01,0.70,0.94\}$. For each case, the first $24$ radial functions have been computed using canonical confluent Heun polynomials derived from the s=2s=-2, $(\ell,m,n)=(2,2,0)$ QNM frequencies; truncation effects have not been included. Different values of $p$ and $q$ correspond to different values of $n^{\prime}$ as discussed in Sec. III. Mixing matrices have been computed according to Eq. (100).
@@
Further, due to the self-adjoint nature of the radial operator with respect to the scalar product (Eq. 13), the space of radial functions is orthogonal London (2023a); Lax (2002); Harvey (1955). Since the canonical confluent Heun polynomial are an orthonormal basis, scalar products between different radial functions, $|{f_{p}}\rangle$ and $|{f_{q}}\rangle$, are equivalent to the simple vector dot product between their vector representations,
@@
Note that, in Eq. (99), canonical confluent Heun polynomials constitute spin-dependent bases. Note also that the radial functions may be normalized in the standard way, $|{f_{p}}\rangle\leftarrow|{f_{p}}\rangle/\sqrt{{\langle{f_{p}}\,|\,{f_{p}}\rangle}}$. In other words, when computed for a single frequency parameter, the radial functions define an orthonormal basis.
A important question in single BH perturbation theory is whether QNMs for one physical case may be used to efficiently represent QNMs in another. For example, it is common to represent the Kerr QNMs’ angular functions in the basis of those for Schwarzschild London (2023a, b); Berti and Klein (2014); O’Sullivan and Hughes (2014); Cook and Zalutskiy (2014); Press and Teukolsky (1973). Here, we are concerned with whether the same is true for the QNMs’ radial functions. For simplicity, we will consider different BH spins, a single QNM frequency for each, and related families of single-frequency solutions to the radial equation. In effect, we ask whether there exists a mixing matrix between sets of radial functions derived from different BH spins [^12] Note that the invertibility of such a matrix was found to be a prerequisite to the existence of the adjoint-spheroidal harmonics presented in Ref. London (2023b)..
To answer this question, it cannot be avoided that, even for a single QNM, different values of BH spin correspond to different QNM frequencies London and Fauchon-Jones (2019); Leaver (1985); Berti and Klein (2014); Berti et al. (2016). When combined with the fact that all forms of Teukolsky’s radial equation (Eq. 10) are frequency dependent, this means that the QNM radial functions for different BH spins are members of different scalar product spaces. While a complete treatment of the question may require the use of a frequency-independent weight function, we remain free to apply any scalar product to the problem. Any choice of scalar product will impact the symmetry properties of $\mathcal{L}_{\xi}$ and the orthogonality properties of the radial functions.
Let the $p^{\text{th}}$ radial function for a given BH spin, $a$, be denoted $\vec{f}_{p}^{\,(a)}$. As mentioned previously, if $\vec{f}_{p}^{\,(a)}$ are defined with QNM frequency $\tilde{\omega}_{{\ell mn}}$, then $\vec{f}_{n}^{\,(a)}$ is the QNM’s radial function. Since each $a$-dependent canonical confluent Heun polynomial basis is orthonormal, Eq. (99) may be interpreted to mean that the scalar product for all $\vec{f}_{p}^{\,(a)}$ is given by the right hand side of Eq. (99c). Since the right hand side of Eq. (99c) has no weight function, we may superficially conclude that the standard vector dot product is a frequency-independent scalar product of potential use.